OpenStudy (anonymous):

7 3 --- > - ---- 8x 16 first part is 7/8x, the sign has a line under it and the second part is -3/16

OpenStudy (anonymous):

\[\frac{7}{8x}\geq \frac{3}{16}\] or \[\frac{7}{8}x\geq \frac{3}{16}\] ??

OpenStudy (anonymous):

second one, but the 3/16 is negative

OpenStudy (anonymous):

ok then it is much easier of the two multiply both sides by 8, divide both sides by 7

OpenStudy (anonymous):

ok let me figure this out...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

1/8 > 3/14

OpenStudy (anonymous):

is what I came up with multiplying both sides by 8 and then dividing by 7

OpenStudy (anonymous):

where is your variable?

OpenStudy (anonymous):

1/8x > 3/14

OpenStudy (anonymous):

hold on, i think you are confused a bit what do you get when you multiply both sides of \[\frac{7}{8}x\geq -\frac{3}{16}\] by \(8\) ?

OpenStudy (anonymous):

when I multiplied 7/8 x by 8 I got 0.875